OpenStudy (king):

Help................find all pairs of integers such that x^2-y^2=104 Help................find all pairs of integers such that x^2-y^2=104 @Mathematics

OpenStudy (anonymous):

\[(x+y)(x-y) = 104\]

OpenStudy (king):

yeah then...

OpenStudy (anonymous):

factorize 104

OpenStudy (king):

yeah i get 52,26 but then i have to find all pairs of integers so is there any other short mthod?

OpenStudy (king):

x=52,y=26

OpenStudy (mathmagician):

104=2*2*2*13. then you have to solve systems of equations: x-y=2 and x+y=2*2*13; x-y=2*2 and x+y=2*13... then you will find all possible pairs